# On the reconstruction limits of complex networks

### v2

@@ Abstract
Network reconstruction consists in retrieving the hidden interaction structure of a system from observations. Many reconstruction algorithms have been proposed, although less research has been devoted to describe their theoretical limitations. In this work, we adopt an information-theoretic perspective and define the reconstructability: The fraction of structural information recoverable from data. The reconstructability depends on the true data generating (TDG) model which is shown to set the reconstruction limit: any algorithm can perform, on average, at best like the TDG model. We show that the reconstructability is related to various performance measures, such as the probability of error and the Jaccard similarity. In an empirical context where the TDG model is unknown, we introduce the reconstruction index as an approximation of the reconstructability. We find that performing model selection is crucial for the validity of the reconstruction index as a proxy of the reconstructability of empirical time series and networks.
@@ I Introduction
Complex systems, such as the brain, are naturally represented by complex networks that encapsulate intricate interactions between neurons or brain regions Bassett and Sporns (2017); Bassett et al. (2018); Lynn and Bassett (2019); Sporns (2013). Network representation unlocks a variety of tools with the potential to unravel not only brain functions and diseases Supekar et al. (2008); Hlinka and Coombes (2012); Forrester et al. (2020), but also gene expressions Wang et al. (2006), epidemics Peixoto (2019); Prasse et al. (2020) and financial distress Musmeci et al. (2013). The problem is that such network representations are seldom measurable experimentally. For example, the collected data are often indirect observations of the interactions, taking the form of counts of interactions or times series. Moreover, these data are noisy, thereby making the network reconstruction task even more intricate Newman (2018); Peixoto (2018); Young et al. (2020); Peel et al. (2022).
@@
The structure-function relationship has recently been investigated in a broader context for random networked processes Murphy et al. (2024) using information-theoretic tools. The study uncovered a duality between reconstructability and predictability, showing that in certain parameter ranges, an increase in predictability corresponds to a decrease in reconstructability, and vice versa. These theoretical findings provide a unifying framework for addressing two a priori unrelated questions : To what extent is a network reconstructible, and how descriptive are the networks we observe in—and reconstruct from—nature?
In this work, we focus on the question of the reconstructability of networks. In Sec. II.1, we describe more precisely the network reconstruction problem using a Bayesian perspective. We identify the crucial role that data generation plays in our ability to reconstruct networks, and link it to the reconstruction limit. Section II.2 formalizes this limit in information-theoretic terms. We find that the mutual information, between the graph and the data used to reconstruct, is the limiting factor of network reconstruction. Section II.3 generalizes the mutual information to a realistic, data-driven scenario—where only the observations are known. As in Ref. Peel et al. (2022), we find that the quality of the model to describe the observations is critical for our ability to describe the reconstruction limit. Finally, in Sec. II.4 we apply our framework to time series and empirical networks.
@@
Fig. 1: Illustration of the network reconstruction context from (a, b) a theoretical perspective and (c, d) an empirical perspective. Panel (a) sketches how the true data generating model (TDG) $M^{*}$ operates, first by generating a graph, then by encoding it into the observations, and finally using these to decode—or reconstruct—the graph. The thickness of the contour line around each graph and data example indicates the probabilities $P(G^{*})$ (top and bottom layers) and $P(X^{*})$ (middle layer). The thickness of the edges connecting the graphs to the data illustrate the likelihood of the TDG $P(X^{*}|G^{*})$, and those connecting the data to a reconstructed graphs, some distribution $P(\hat{G}|X^{*})$. In panel (b), we illustrate in red the reconstructible information, utilizing an information-theoretic perspective. This information is part of the total information of $G^{*}$ and $X^{*}$—in blue and orange, respectively—and is also a fraction of the partial information of $G^{*}$ needed to completely reconstruct it (blue and red). Panels (c, d) show the analog of (a, b) when model $M^{*}$ is unknown, where in panel (c) a single datum is accessible and reconstruction is done by a candidate model $M$, a priori different from $M^{*}$. In panel (d), we illustrate how $M$ and $M^{*}$ may overlap in the information they reconstruct—the information intersection (i.e., the correctly recovered information) and difference (i.e., the missing or spurious information). The reconstructability $U(G^{*}|X^{*})$ and the reconstruction index $\Psi_{M}$ are defined in subsection II.2.2 and subsection II.3.1, respectively.
II Results
II.1 Network reconstruction
@@
We formulate the network reconstruction problem as follows. Let $g^{*}\in\mathcal{G}$ be some graph of $N$ nodes that represents the structure of the interactions between each pair of components in a system, where $\mathcal{G}$ is the set of all graphs of $N$ nodes. The graph may be directed and weighted Peixoto (2024), but we restrict our discussion to undirected and unweighted, for simplicity. This graph structure is a priori unknown to us, although it is indirectly observed through some data, denoted $x^{*}$, which may take any value in the set $\mathcal{X}$. This data can take many forms—time series, pairwise measurements, etc.—and we assume it to be generated using $g^{*}$. The goal of network reconstruction is to infer the graph $g^{*}$ from the data $x^{*}$.
@@
II.1.1 Data generation process
Fig. 2: Performance comparison between the TDG model and heuristic reconstruction algorithms. In both panels, we show the area under the receiver operating characteristic curve (AUC) of the reconstruction models as a function of a parameter of the dynamics that generated the data: (a) the Susceptible-Infection-Susceptible dynamics and (b) Glauber dynamics (see Table 2 for the definitions of the dynamics). We generated graphs of $N=100$ nodes with the Erdős-Rényi model (Eq. (4)), where the number of edges is $E=250$. We also generated time series of $T=500$ time steps; the parameters other than the infection probability $\lambda$ and the coupling constant $J$ (which are fixed within the likelihood during the inference of the TDG) are specified in Table 2. The curves represent averages over 24 realizations, and the shaded regions around each curve indicate 90% confident intervals from the mean.
A Bayesian reconstruction model, composed of the two random variables $G$ and $X$, reflects our assumptions about how the unobserved graph and observed data came to be. In other words, the model $M=(G,X)$ represents a generative process for the pairs $(g^{*},x^{*})$. Accordingly, there are many reconstruction models that may describe the data to various degrees of correctness. Throughout this work, we assume the existence of a unique generative process, referred to as the true data-generating (TDG) model $M^{*}=(G^{*},X^{*})$, which truly produced the graph $g^{*}$ and the observed data $x^{*}$ with probabilities $P(G^{*}=g^{*})$ and $P(X^{*}=x^{*}|G^{*}=g^{*})$, respectively. In turn, any reconstruction model may be described by a reconstructed random graph $\hat{G}$, that depends on $X^{*}$. The complete process consisting of the graph and data generation followed by the reconstruction of the graph is therefore described by the random variable triplet $(G^{*},X^{*},\hat{G})$, whose joint probability distribution is
@@ II.1.1 Data generation process
for all $g\in\mathcal{G}$, such that $P(G|X)$ is given by Eq. (1). Note that the reconstructed random graph $\hat{G}$ and the random graph $G$ of model $M$ conceptually describe two different quantities, although they are related through Eq. (3). Indeed, $\hat{G}$ appears in the reconstruction process involving the TDG and $G$ is part of a completely separate generative process. In other words, $\hat{G}$ depends explicitly on $M^{*}$, via $P(\hat{G}|X^{*})$, whereas $M$ is independent from it (i.e., $P(G,X|G^{*},X^{*})=P(G,X)$). The consideration that $\hat{G}$ is, in fact, resulting from a Bayesian procedure through a generative model $M$, instead of any—potentially nongenerative—model, will prove useful in the following sections.
From an information-theoretic perspective, data generation encodes information about the graph $G^{*}$ into potentially noisy observations $X^{*}$, while network reconstruction decodes these observations back into a graph $\hat{G}$ as shown in Fig. 1(a–b). The encoding of $G^{*}$ into $X^{*}$ is generally lossy, meaning that only a fraction of its information can be recovered; the rest being lost in the process. In turn, any reconstruction model $M$ different in distribution from $M^{*}$ may therefore recover a fraction of the reconstructible information while potentially introducing spurious information through their inductive biases [see Fig. 1(c–d)], resulting in a degradation of performance. This is well shown in Fig. 2 through reconstruction performance, where we compare the area under the receiver operating characteristic curve (AUC) of the TDG reconstruction model with the AUC of three heuristic reconstruction algorithms. In this example, the random graph of the TDG model is a simple Erdős-Rényi (ER) graph, whose graph prior is
@@
with $E$ being the (given) number of edges in the graph (see Appendix IV.1 for details). Then, the data consists in time series of binary node activities, generated using the Glauber dynamics model Glauber (1963) in panel (a) and the Susceptible-Infected-Susceptible (SIS) dynamics model Pastor-Satorras et al. (2015) in panel (b) (see Appendix IV.2 for details). As a comparison, we use three different well-known reconstruction algorithms: the correlation matrix method Kramer et al. (2009), Granger causality method Schreiber (2000) and the transfer entropy method Seth (2005) (see Appendix IV.3 for details). The results in Fig. 2 show quite unambiguously and unsurprisingly that the TDG model outperforms the reconstruction heuristics.
@@
Yet, even the TDG reconstruction model cannot reconstruct the graph perfectly. For instance, in Fig 2(b), the AUC of the TDG model tends to $\frac{1}{2}$—equivalent to random guessing—when the coupling also goes to zero. In this scenario, $X^{*}$ and $G^{*}$ are independent and it is actually impossible to reconstruct the graph, since any graph could have generated the data with the exact same probability. The same phenomenon occurs to a lesser extent for the other coupling values as well as for the SIS dynamics, where the TDG model performance is imperfect for every infection probability. These imperfections are attributed to the lost information in the encoding of $G^{*}$; no model can extract more information than what is contained in $X$. In practice, the encoding’s loss stems from many sources, for example noise in the dynamics and degeneracy, where many networks lead to similar dynamics. The degeneracy phenomenon is well-established in computational neuroscience Prinz et al. (2004); Cropper et al. (2016) and has more recently appeared in network science Prasse and Mieghem (2022) too. A reconstruction limit independent of the reconstruction algorithm clearly exists, where a perfect reconstruction is simply not attainable even in the best-case scenario. This is a key insight that we will explore in the following sections (especially Sec. II.3).
II.1.2 Reconstructing a single edge
@@
To gain better intuition about this reconstruction limit, we consider the reconstruction of a graph that may only contain a single edge. Let $G^{*}$ be a random graph of two nodes, that may be connected by a single edge with probability $p$, and disconnected with probability $1-p$. This edge is observed through a noisy process $X^{*}=(X_{1},...,X_{T})$ with $T$ time steps, where $X_{i}$ is a binary variable that takes the value $1$ if the edge has been observed and 00 otherwise. We assume that the noisy process can induce true positives and false positives, each with known probabilities $q$ and $r$, respectively, making the reconstruction problem more challenging. The likelihood $P(n|a)$ that the edge has been observed $n$ times, given that it is present ($a=1$) or not ($a=0$), is a binomial distribution:
@@
Figure 3 shows the behavior of $P(a=1|n)$ when varying the number $n$ of times the edge is observed and the true positive probability $q$. Assuming that $r<q$, observed edges are mostly true positives and thus the edge is predicted to exist if $n$ is sufficiently large; otherwise, it is not since the expected number of true and false positives don’t match the observations. Conversely, if $r>q$, then most observed edges are false positives, meaning that $g^{*}$ is more likely to contain an edge when $n$ is small. Interestingly, the edge becomes more challenging to reconstruct as $q$ gets closer to $r$, where the probability to reconstruct the edge approaches $\frac{1}{2}$ (see Fig. 3(a)). In this regime, $a$ and $1-a$ are interchangeable and it becomes impossible to tell if the edge exists or not—any attempt at reconstructing this graph would be unfruitful. This is precisely the intuition we want to capture with the reconstruction limit: When is there enough information to properly reconstruct the structure, or to what extent is a system’s structure reconstructible? In the next section, we present an information-theoretic framework that quantifies this limit.
II.2 Information-theoretic reconstruction limits
As discussed above, we can think of the TDG process $X^{*}$ as a noisy encoding of the true graph $G^{*}$. This amount of encrypted information is fundamentally limiting our ability to reconstruct $G^{*}$ accurately; it is impossible to recover more information than what is contained in the data. This also means that the limit is independent of the reconstruction models or algorithms. Any reconstruction algorithm therefore aims to extract as much of the encrypted information as possible, some being more efficient than others.
II.2.1 Entropy
@@
×$\qquad\qquad\qquad\times\log P(G=g|X=x)\,.$ (10)
@@ II.2.2 Network reconstructability
$H(G^{*}|\hat{G})=-\mathbb{E}_{G^{*},\hat{G}}\quantity[\log P(G^{*}|\hat{G})]\,$ (12)
is the entropy of the true graph given the reconstructed one. The conditional probability $P(G^{*}|\hat{G})=P(G^{*},\hat{G})/P(\hat{G})$ is such that both $P(G^{*},\hat{G})$ and $P(\hat{G})$ are marginal distributions of $P(G^{*},X^{*},\hat{G})$ [Eq. (2)]. Three observations regarding this performance measure are in order. First, the quantity $I(G^{*};\hat{G})$ may be interpreted as measuring the similarity between the information contents of $G^{*}$ and $\hat{G}$. The higher it is, the more similar $G^{*}$ and $\hat{G}$ are and the better is the reconstruction. Conversely, when $I(G^{*};\hat{G})=0$, it is minimized and both graphs are independent from one another. Note that similar mutual information measures have been used as a performance measure in the context of community detection for comparing pairs of partitions Newman et al. (2020); Jerdee et al. (2024).
Second, $I(G^{*};\hat{G})$ is related to the probability of error, defined as
@@
This lower bound on the probability of error is minimized when $I(G^{*};\hat{G})$ is maximized.
@@
is the mutual information between the true graph $G^{*}$ and the data process $X^{*}$. Intuitively, $I(G^{*};X^{*})$ quantifies the amount of reconstructible information that both $X^{*}$ and $G^{*}$ share—i.e., the amount of information that $X^{*}$ contains about $G^{*}$ [see Fig. 1(b)]. The mutual information $I(G^{*};X^{*})$ also sets the maximum in reconstruction performance as measured by $I(G^{*};\hat{G})$: It is the reconstruction limit.
The mutual information $I(G^{*};X^{*})$ is itself bounded between 00 and $H(G^{*})$ Cover and Thomas (2006). When $I(G^{*};X^{*})=0$, $X^{*}$ and $G^{*}$ are independent and thus the data $X^{*}$ contains no information about the graph $G^{*}$. In turn, it is impossible for any reconstruction model $M$ to extract information from the data, regardless of its specification. When $I(G^{*};X^{*})=H(G^{*})$, the data $X^{*}$ contains all the information about the graph $G^{*}$. In this case, it is in principle possible to perfectly reconstruct the graph without any error, assuming the model $M$ is optimal.
Since the value of $I(G^{*};X^{*})$ depends on the amount of information $H(G^{*})$ that needs to be extracted, it is easier to reason about it in terms of proportions. Thus, we define the reconstructability $U(G^{*}|X^{*})$ of $G^{*}$ from $X^{*}$ as the uncertainty coefficient
$U(G^{*}|X^{*})=\frac{I(G^{*};X^{*})}{H(G^{*})}\,.$ (18)
The reconstructability has been described thoroughly in Ref. Murphy et al. (2024) and helped unveiling a special duality between our ability to predict the time evolution of a system and our ability to reconstruct the interactions between its constituents. As it is a normalized version of the mutual information upper bound $I(G^{*};X^{*})$, the reconstructability is bounded between 00 and $1$. When $U(G^{*}|X^{*})=0$, any attempt at reconstruction is futile, whereas it is theoretically possible to decode all the information when $U(G^{*}|X^{*})=1$. As such, the reconstructability is a measure of the average proportion of information that can be extracted from the data about the graph. For instance, when it is equal to $\frac{1}{2}$, it precisely means that half of the graph information is, on average, contained in the data and that in turn half of it can possibly be reconstructed. We stress that $U(G^{*}|X^{*})=\frac{1}{2}$ may not be directly interpreted as half of the graph’s edges being reconstructible. Rather, information may generally be distributed in a heterogeneous way over the graph’s structure, as a single bit of information may reconstruct more than one edge in the graph depending on how correlated they are.
Going back to the earlier example of a single edge of Sec. II.1.2, we can perform the complete calculation analytically. First, we can calculate the entropy of the prior,
@@
recalling that $h(p)$ is the binary entropy defined above, and the entropy of the posterior is
@@
×$\qquad\qquad\times h\left(\frac{p}{p+\eta^{T-n}\lambda^{n}(1-p)}\right)\,,$ (20)
This leads to the reconstructability of the edge using Eq. (18), which is plotted in Fig. 3. As expected, the reconstructability typically increases as the number of observations $T$ increases, even reaching 1 in some cases, e.g., when $q\to 1$. Also, notice how the reconstructability is zero for every value of $T$ when the true positive probability $q$ is equal to false positive probability $r$. This shows that as true positives and false positives become indistinguishable, the edge becomes impossible to reconstruct.
II.2.3 Optimal reconstruction performance
The reconstruction limit corresponds to the maximum performance, as measured by $I(G^{*};\hat{G})$, achievable by any algorithm. Hence, any reconstruction model that is capable of reaching this limit, i.e., $I(G^{*};\hat{G})=I(G^{*};X^{*})$, is said optimal in its reconstruction abilities. It is not surprising that the TDG model is optimal according to this definition. As a result, the reconstructability $U(G^{*}|X^{*})$ can also be interpreted as a reconstruction performance measure of the TDG model $M^{*}$.

### v3

@@ Abstract
Network reconstruction consists in retrieving the hidden interaction structure of a system from observations. Many reconstruction algorithms have been proposed, although less research has been devoted to describe their theoretical limitations. In this work, we take a first-principles approach and build on our earlier definition of reconstructability—the fraction of structural information recoverable from data. We relate this quantity to the true data-generating (TDG) process and delineate an information-theoretic reconstruction limit, i.e., the upper bound of the mutual information between the true underlying graph and any graph reconstructed from observations. These concepts lead us to a principled numerical method to assess the validity of empirically reconstructed networks, based on model selection and a quantity we introduce: the reconstruction index. This index approximates the reconstructability from data, quantifies the variability of the reconstructed network ensemble, and is shown to predict reconstruction error without requiring knowledge of the true underlying network. We characterize this method and test it on empirical time series and networks.
@@ I Introduction
Complex systems, such as the brain, are naturally represented by complex networks that encapsulate intricate interactions between neurons or brain regions Bassett and Sporns (2017); Bassett et al. (2018); Lynn and Bassett (2019); Sporns (2013). Network representation unlocks a variety of tools with the potential to unravel not only brain functions and diseases Supekar et al. (2008); Hlinka and Coombes (2012); Forrester et al. (2020), but also gene expressions Wang et al. (2006), epidemics Peixoto (2019); Prasse et al. (2020) and the propagation of financial distress Musmeci et al. (2013). The main challenge is that such network representations are seldom measurable experimentally. For example, the collected data are often indirect observations of the interactions, taking the form of counts of interactions or times series. Moreover, these data are noisy, thereby making the network reconstruction task even more intricate Newman (2018); Peixoto (2018); Young et al. (2020); Peel et al. (2022).
@@
An information-theoretic description of random networked processes has recently led to a broader understanding of this so-called structure-function relationship, linking predictability to reconstructability in complex networks through mutual information Murphy et al. (2024). This description revealed a duality between reconstructability and predictability, showing that in certain parameter ranges, an increase in predictability corresponds to a decrease in reconstructability, and vice versa. In this work, we aim to further this theory on the network reconstruction front, providing an information-theoretic bedrock to such applications.
In the network reconstruction problem, our task is to infer a graph likely to have generated some observed data. We first formally present this problem and the related mathematical concepts in Sec. II. Then, we revisit and adapt the framework of Ref. Murphy et al. (2024) in Sec. III, allowing us to interpret the reconstruction problem in information-theoretic terms. In doing so, we demonstrate the existence of an algorithm‑independent limit to network reconstruction—the reconstructability $\Psi^{*}$ [see Fig. 1(a,b)]—which bounds from above the mutual information between the true underlying network and the reconstructed one. Inspired by this limit, we present and characterize in Sec. IV a principled numerical method to assess the validity of reconstructed networks in an empirical setting (i.e., hidden generative process, one or few observations). Our method is based on the reconstruction index, denoted $\psi_{M}$ for some reconstruction model $M$, which is an approximation of the reconstructability $\Psi^{*}$ that measures the dispersion of the reconstructed graph ensemble. The reconstruction index is shown to predict the reconstruction error without knowing the true underlying graph [see Fig. 1(c,d)], assuming our modeling assumptions are aligned with the true underlying process. Finally, we apply our method to real systems in Sec. V.
@@
Fig. 1: Illustration of the network reconstruction context from (a, b) a theoretical perspective and (c, d) an empirical perspective. Panel (a) sketches how the true data generating model (TDG) $M^{*}$ operates, first by generating a graph, then by encoding it into the observations, and finally using these to decode—or reconstruct—the graph. The thickness of the contour line around each graph and data example indicates the probabilities $P(G^{*})$ (top and bottom layers) and $P(X^{*})$ (middle layer). The thickness of the edges connecting the graphs to the data illustrate the likelihood of the TDG $P(X^{*}|G^{*})$, and those connecting the data to a reconstructed graphs, some distribution $P(\hat{G}|X^{*})$. In panel (b), we illustrate in red the reconstructible information, utilizing an information-theoretic perspective. This information is part of the total information of $G^{*}$ and $X^{*}$—in blue and orange, respectively—and is also a fraction of the partial information of $G^{*}$ needed to completely reconstruct it (blue and red). Panels (c, d) show the analog of (a, b) when the model $M^{*}$ is unknown, where in panel (c) a single datum is accessible and reconstruction is done by a candidate model $M$, a priori different from $M^{*}$. In panel (d), we illustrate how $M$ and $M^{*}$ may overlap in the information they reconstruct—the information intersection (i.e., the correctly recovered information) and difference (i.e., the missing or spurious information). The reconstructability $\Psi^{*}$ and the reconstruction index $\psi_{M}$ are defined in subsection III.2 and subsection IV.1, respectively.
II Network reconstruction
@@
We formulate the network reconstruction problem following the illustration in Fig. 1(c). Let $g^{*}\in\mathcal{G}$ be some graph of $N$ nodes that represents the structure of the interactions between each pair of components in a system, where $\mathcal{G}$ is the set of all graphs of $N$ nodes. The graph may be directed and weighted Peixoto (2024), but we restrict our discussion to undirected and unweighted, for simplicity. This graph structure is a priori unknown to us, although it is indirectly observed through some data, denoted $x^{*}$, which may take any value in the set $\mathcal{X}$. This data can take many forms—time series, pairwise measurements, etc.—and we assume it to be generated using $g^{*}$. In what follows, we will further assume that $x^{*}$ is in fact a $N\times T$ matrix corresponding to $N$ coupled time series of length $T$, but we stress that our analysis may apply to any type of networked data. The goal of network reconstruction is to infer the graph $g^{*}$ from the data $x^{*}$.
@@
II.1 Data generation process
Fig. 2: Performance comparison between the TDG model and heuristic reconstruction algorithms. In both panels, we show the area under the receiver operating characteristic curve (AUC) of the reconstruction models as a function of a parameter of the model that generated the data: (a) the Susceptible-Infection-Susceptible (SIS) dynamics and (b) Glauber dynamics (see Table 2 for the definitions of the dynamics). We generated graphs of $N=100$ nodes with the Erdős-Rényi model (Eq. (4)), where the number of edges is $E=250$. We also generated time series of $T=500$ time steps; the parameters other than the infection probability $\lambda$ and the coupling constant $J$ (which are fixed within the likelihood during the inference of the TDG) are specified in Table 2. Each data point corresponds to the AUC average over 24 reconstruction experiments, each experiment with different realizations of $G^{*}$ and $X^{*}$, and the shaded regions around the points show a 90% confident interval from the mean. For further technical details, see Sec. II.1.
[…]
@@ II.1 Data generation process
for all $g\in\mathcal{G}$, such that $P(G|X)$ is given by Eq. (1). Note that the reconstructed random graph $\hat{G}$ and the random graph $G$ of model $M$ conceptually describe two different quantities, although they are related through Eq. (3). Indeed, $\hat{G}$ appears in the reconstruction process involving the TDG and $G$ is part of a completely separate generative process. In other words, $\hat{G}$ depends explicitly on $M^{*}$, via $P(\hat{G}|X^{*})$, whereas $M$ is independent from it (i.e., $P(G,X|G^{*},X^{*})=P(G,X)$). The consideration that $\hat{G}$ is, in fact, resulting from a Bayesian procedure through a generative model $M$, instead of any—potentially nongenerative—algorithm such as the inverse correlation method Kramer et al. (2009), will prove useful in the following sections.
From an information-theoretic perspective, data generation encodes information about the graph $G^{*}$ into potentially noisy observations $X^{*}$, while network reconstruction decodes these observations back into a graph $\hat{G}$ as shown in Fig. 1(a, b). The encoding of $G^{*}$ into $X^{*}$ is generally lossy, meaning that only a fraction of its information can be recovered; the rest being lost in the process. In turn, any reconstruction model $M$ different in distribution from $M^{*}$ may therefore recover a fraction of the reconstructible information while potentially introducing spurious information through their inductive biases [see Fig. 1(c,d)], resulting in a degradation of performance.
This is well shown in Fig. 2 through reconstruction performance, where we used two synthetic TDG processes to compare the TDG reconstruction model performance with that of three heuristic reconstruction algorithms. In this experiment, we sample the true graph with probability
@@
which corresponds to the Erdős-Rényi (ER) model, with $E$ being the (given) number of edges in the graph; and sample time series from the Susceptible-Infected-Susceptible (SIS) model in panel (a) and Glauber models in panel (b). For further details regarding the graph and data models, see Appendices A and B. Then, assuming a given reconstruction model, we compare the reconstructed graph with original one using the area under the receiver operating characteristic curve (AUC) to measure reconstruction performance. We repeat this set up for many parameter values to populate the AUC performance curves in Fig. 2. As a comparison, we use three different well-known reconstruction algorithms: the correlation matrix method Kramer et al. (2009), Granger causality method Schreiber (2000) and the transfer entropy method Seth (2005) (see Appendix C for details). The results in Fig. 2 show quite unambiguously and unsurprisingly that the TDG model outperforms the reconstruction heuristics.
@@
Yet, even the TDG reconstruction model cannot reconstruct the graph perfectly. For instance, in Fig 2(b), the AUC of the TDG model tends to $\frac{1}{2}$—equivalent to random guessing—when the coupling also goes to zero. In this scenario, $X^{*}$ and $G^{*}$ are independent and it is actually impossible to reconstruct the graph, since any graph could have generated the data with the exact same probability. The same phenomenon occurs to a lesser extent for the other coupling values as well as for the SIS dynamics, where the TDG model performance is imperfect for every infection probability. These imperfections are attributed to the lost information in the encoding of $G^{*}$; no model can extract more information than what is contained in $X$. In practice, the encoding’s loss stems from many sources, for example noise in the dynamics and degeneracy, where many networks lead to similar dynamics. The degeneracy phenomenon is well-established in computational neuroscience Prinz et al. (2004); Cropper et al. (2016) and has more recently appeared in network science Prasse and Mieghem (2022) too. A reconstruction limit independent of the reconstruction algorithm clearly exists, where a perfect reconstruction is simply not attainable even in the best-case scenario. This is a key insight that we will explore in the following sections (especially Sec. IV).
II.2 Reconstructing a single edge
@@
To gain better intuition about this reconstruction limit, we consider the reconstruction of a graph that may only contain a single edge. Let $G^{*}$ be a random graph of two nodes, that may be connected by a single edge with probability $p$, and disconnected with probability 1p1-p. This edge is observed through a noisy process $X^{*}=(X_{1},...,X_{T})$ with $T$ time steps, where $X_{i}$ is a binary variable that takes the value $1$ if the edge has been observed and 0 otherwise. We assume that the noisy process can induce true positives and false positives, each with known probabilities $q$ and $r$, respectively, making the reconstruction problem more challenging. The likelihood $P(n|a)$ that the edge has been observed $n$ times, given that it is present ($a=1$) or not ($a=0$), is a binomial distribution:
@@
Figure 3 shows the behavior of $P(a=1|n)$ when varying the number $n$ of times the edge is observed and the true positive probability $q$. Assuming that $r<q$, observed edges are mostly true positives and thus the edge is predicted to exist if $n$ is sufficiently large; otherwise, it is not since the expected number of true and false positives don’t match the observations. Conversely, if $r>q$, then most observed edges are false positives, meaning that $g^{*}$ is more likely to contain an edge when $n$ is small. Interestingly, the edge becomes more challenging to reconstruct as $q$ gets closer to $r$, where the probability to reconstruct the edge approaches $\frac{1}{2}$ (see Fig. 3(a)). In this regime, $a$ and 1a1-a are interchangeable and it becomes impossible to tell if the edge exists or not—any attempt at reconstructing this graph would be unfruitful. This is precisely the intuition we want to capture with the reconstruction limit: When is there enough information to properly reconstruct the structure, or to what extent is a system’s structure reconstructible? In the next section, we present an information-theoretic framework that quantifies this limit.
III Information-theoretic reconstruction limits
As discussed above, we can think of the TDG process $X^{*}$ as a noisy encoding of the true graph $G^{*}$. This amount of encoded information is fundamentally limiting our ability to reconstruct $G^{*}$ accurately; it is impossible to recover more information than what is contained in the data. This also means that the limit is independent of the reconstruction models or algorithms. Any reconstruction algorithm therefore aims to extract as much of the encoded information as possible, some being more efficient than others.
III.1 Entropy
@@
$\qquad\qquad\qquad\times\log P(G=g|X=x)\,.$ (10)
@@ III.2 Network reconstructability
[…]
Second, I(G;G^)I($^{*};\hat{G}$) is related to the probability of error, defined as
@@
This lower bound on the probability of error is minimized when I(G;G^)I($^{*};\hat{G}$) is maximized.
@@
is the mutual information between the true graph $G^{*}$ and the data process $X^{*}$. Intuitively, $I(G^{*};X^{*})$ quantifies the amount of reconstructible information that both $X^{*}$ and $G^{*}$ share—i.e., the amount of information that $X^{*}$ contains about $G^{*}$ [see Fig. 1(b)]. The mutual information $I(G^{*};X^{*})$ also sets the maximum in reconstruction performance as measured by I(G;G^)I($^{*};\hat{G}$): It is the reconstruction limit.
The mutual information $I(G^{*};X^{*})$ is itself bounded between 0 and $H(G^{*})$ Cover and Thomas (2006). When $I(G^{*};X^{*})=0$, $X^{*}$ and $G^{*}$ are independent and thus the data $X^{*}$ contains no information about the graph $G^{*}$. In turn, it is impossible for any reconstruction model $M$ to extract information from the data, regardless of its specification. When $I(G^{*};X^{*})=H(G^{*})$, the data $X^{*}$ contains all the information about the graph $G^{*}$. In this case, it is in principle possible to perfectly reconstruct the graph without any error, assuming the model $M$ is optimal, i.e., it can extract all the available information. We will further explore this notion of optimality in Sec. III.3.
Since the value of $I(G^{*};X^{*})$ depends on the amount of information $H(G^{*})$ that needs to be extracted, it is easier to reason about it in terms of proportions. Thus, we define the reconstructability $\Psi^{*}$ of $G^{*}$ from $X^{*}$ as the uncertainty coefficient
$\Psi^{*}=\frac{I(G^{*};X^{*})}{H(G^{*})}\,.$ (18)
The reconstructability has been described thoroughly in Ref. Murphy et al. (2024) and helped unveiling a special duality between our ability to predict the time evolution of a system and our ability to reconstruct the interactions between its constituents. As it is a normalized version of the mutual information upper bound $I(G^{*};X^{*})$, the reconstructability is bounded between 0 and $1$. When $\Psi^{*}=0$, any attempt at reconstruction is futile, whereas it is theoretically possible to decode all the information when $\Psi^{*}=1$. As such, the reconstructability is a measure of the average proportion of information that can be extracted from the data about the graph. For instance, when it is equal to $\frac{1}{2}$, it precisely means that half of the graph information is, on average, contained in the data and that in turn half of it can possibly be reconstructed. We stress that $\Psi^{*}=\frac{1}{2}$ may not be directly interpreted as half of the graph’s edges being reconstructible. Rather, information may generally be distributed in a heterogeneous way over the graph’s structure, as a single bit of information may reconstruct more than one edge in the graph depending on how correlated they are.
Going back to the earlier example of a single edge of Sec. II.2, we can perform the complete calculation analytically. First, we can calculate the entropy of the prior,
@@
recalling that $h(p)$ is the binary entropy defined below Eq. (14), and the entropy of the posterior is
@@
$\qquad\qquad\times h\left(\frac{p}{p+\eta^{T-n}\lambda^{n}(1-p)}\right)\,.$ (20)
This leads to the reconstructability of the edge using Eq. (18), which is plotted in Fig. 3(b). As expected, the reconstructability typically increases as the number of observations $T$ increases, even reaching 1 in some cases (e.g., when $q\to 1$). Also, notice how the reconstructability is zero for every value of $T$ when the true positive probability $q$ is equal to false positive probability $r$. This shows that as true positives and false positives become indistinguishable, the edge becomes impossible to reconstruct.
III.3 Optimal reconstruction performance
The reconstruction limit corresponds to the maximum performance, as measured by I(G;G^)I($^{*};\hat{G}$), achievable by any algorithm. Hence, any reconstruction model that is capable of reaching this limit, i.e., $I(G^{*};\hat{G})=I(G^{*};X^{*})=\Psi^{*}H(G^{*})$, is said optimal in its reconstruction abilities. It is not surprising that the TDG model is optimal according to this definition. As a result, the reconstructability $\Psi^{*}$ can also be interpreted as a reconstruction performance measure of the TDG model $M^{*}$.